If you are traveling at 55 miles per hour, what formula would you use to find the distance you will travel over variuos time intervals?

Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
s = 55 mph
---
d = 55 * t
where:
d =distance in miles
t =time in hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php